OpenStudy (anonymous):

How many liters of methane gas (CH4) need to be combusted to produce 12.4 liters of water vapor, if all measurements are taken at the same temperature and pressure? Show all of the work used to solve this problem. CH4 (g) + 2 O2 (g) ----> CO2 (g) + 2 H2O (g)

OpenStudy (anonymous):

6.2

OpenStudy (cherylim23):

You know moles = volume/molar volume (Vm) n(H2O) = 12.4/(Vm) n(CH4)=n(H2O) * 0.5 = 6.2/Vm V(CH4) = 6.4/Vm *Vm = 6.2L

OpenStudy (anonymous):

No.. This is not a practice exam. This is a module quiz from FLVS and it's cheating if you use this answer. Plus it has a obvious mistake in it. Mr. Carlyle FLVS Chemistry Instructor This question is a violation of the OpenStudy Guidelines. The question is from an online school plus a violation by solicitation for members to unknowingly assist them cheating on an exam. 30653065chem30653065